OpenStudy (anonymous):

Factor the trinomial: x2 - 14x + 24 A. (x - 3)(x - 8) B. (x - 2)(x - 12) C. (x - 4)(x + 12) D. prime

OpenStudy (anonymous):

I get B im not sure tho

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

So we want `factors` of positive 24 that will also `add` to -14. Let's check B a sec. The given factors are -2 and -12. \[\Large -2\times-12=24\]\[\Large -2+-12=-14\] Those seem to be the correct values that we want! :)

OpenStudy (anonymous):

its -12 x -2

OpenStudy (anonymous):

\[x^{2}-12x-2x+24 \]

OpenStudy (anonymous):

\[(x ^{2}-12)(-2x+24)\]

OpenStudy (anonymous):

@zepdrix \[x(x-12)2x(-1+24)?\]

zepdrix (zepdrix):

Oh you're trying to see the factors by grouping? Start by factoring an x out of the first two terms. \[\Large x^{2}-12x-2x+24 \quad=\quad x(x-12)-2x+24\] Then factor a -2 out of each of the second terms.\[\Large x\color{royalblue}{(x-12)}-2\color{royalblue}{(x-12)}\]

OpenStudy (anonymous):

Soo im corecct B

OpenStudy (anonymous):

\[(x-12)(x-2)\]

OpenStudy (anonymous):

1 Question what would it have to be in order for it to be prime ?i don't get prime.

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

oh prime? :o ummmm

OpenStudy (anonymous):

Yeah idk that

zepdrix (zepdrix):

Think about what it means for a `number` to be prime. A number is prime if it's only factors are 1 and itself, right? Example: 7 has factors 1x7, but no other factors. So it is a prime number.

zepdrix (zepdrix):

For a `polynomial` to be prime, it can only be written as the product of 1 and itself.\[\Large x^2-2x+7 \quad=\quad (1)(x^2-2x+7)\]

zepdrix (zepdrix):

In the example above, there is no way for us to write x^2-2x+7 as any other factors. (at least not with real integer values).\[\Large x^2-2x+7 \quad \ne \quad(\qquad)(\qquad)\]

zepdrix (zepdrix):

Maybe I'm making this more complicated than it needs to be :) lol It simply means there is no way for us to factor the polynomial.

OpenStudy (anonymous):

NO no i get it thanks